OpenStudy (anonymous):

You roll a blue die and a yellow die. What is P(the sum of the dice is at least 6 | the blue die shows a 2)? Write fractions using a slash and reduce them to lowest terms.

OpenStudy (anonymous):

@ybarrap

OpenStudy (anonymous):

b/y 1/2

OpenStudy (anonymous):

how did u get that

OpenStudy (anonymous):

there is one blue and one yellow...that is 2. if you roll both its a 50 50 chance so 1/2

OpenStudy (anonymous):

ok thnk u

OpenStudy (anonymous):

yeahhhhh:)

OpenStudy (anonymous):

what abt this?:

OpenStudy (anonymous):

You roll two dice. Let event A be "The first die shows a 1 or a 6" and event B be "The second die shows an even number." What is P(A | B)?

OpenStudy (ybarrap):

In the first problem, if the blue die lands 2 then in order for the sum of the blue and yellow to be 6 or greater, the yellow needs to be a 4, 5 or 6, since 2 + 4, 2 +5 and 2 + 6 are all sums greater than or equal to 6, The chance that yellow is 4,5 or 6 is 3/6 or 1/2.

OpenStudy (anonymous):

its not 1/3??

OpenStudy (anonymous):

its 1/2 sleek-feathered one

OpenStudy (anonymous):

No @MROAKS1432 the answer is 1/3

OpenStudy (anonymous):

how did u get that

OpenStudy (anonymous):

Because if P(A) = P(1 or 6) then: P(1) + P(6) and then you can get your answer: (1/6)+(1 / 6) = 2/6 =1/3

OpenStudy (anonymous):

so it is 1/3. @ybarrap i don't really understand wat u said though

OpenStudy (anonymous):

The correct answer is 1/3. Can I get a medal :)

OpenStudy (anonymous):

it says that i already gave out a medal.. sorry

OpenStudy (anonymous):

:( its ok but glad to help :)

OpenStudy (anonymous):

thanks. can any of u guys open the attachment??? :(

OpenStudy (ybarrap):

My reply above was for your first problem. For the second problem P(A|B) = P(A and B)/P(B), but each are independent, so P(A|B) = P(A)P(B)/P(B) = P(A) = P(1) + P(6) = 1/6 + 1/6 = 1/3 as @AwesomeB stated

OpenStudy (anonymous):

yay :)

OpenStudy (anonymous):

ok what abt the ATTACHMENT above

OpenStudy (anonymous):

ill repost it as a new question & then i'll tag u guys??

OpenStudy (anonymous):

The attachment you said was a lot of random numbers and letters

OpenStudy (ybarrap):

I think that the ..rtf part of your doc name is causing problems, my browser is interpreting as a plain text, so the rtf-encoding is showing not the decoded rtf

OpenStudy (anonymous):

ugh! so there's no way u guys can view it?? :(

OpenStudy (anonymous):

no unless you can change it

OpenStudy (ybarrap):

change the name, remove one of the periods

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

even though I change the name it stays a an "rtf"..